### v1

@@ Appendix D Detailed Data Preparation Procedure
    Add 0/50/100 grams of carbohydrate to the three copies at the beginning of exercise (the 7th last time step),
@@
    Add 0/2.5/5 units of insulin to the three copies throughout exercise (the last 7 time steps),
@@
    Add nothing/50 grams of carb/10 units of insulin to the three copies at the beginning of exercise,
@@
    Change the recorded heart rate in the last 7 time steps to typical medium intensity aerobic (80,90,100,110,120,130,120)/ typical interval (80,170,80,170,80,170,80)/typical high intensity resistance (160,170,180,170,160,180,160).
To make sure the labels are not always the same for the same type of intervention, we then switch the 2nd and the 3rd copy with probability 0.5 (again simulated with the same numpy random number generator). Finally we compute the corresponding class label of the intervention set and add it to the causal task solution data set. At the end of this step, for each exercise instance, one should end up with a 4-stack time series of shape (4,54,5).
Next, we perform a round of data cleaning. We removed exercise instances with 0 or nan in heart rate feature, and replaced nan in other features with 0. After this step, we ended up with 94 exercise instances, forming a data set of size (94,4,54,5) and a causal task solution set of size (94,3).
@@ Appendix E Model Implementations and Hyperparameters
For each model mentioned in the experiment section, here we offer a detailed description of its computation equations and the hyperparameters used. Throughout this section we are given an exercise time series of 54 time steps (corresponding to 54 5-min intervals that made up the time window starting from 4 hours before exercise and ending at 30 minutes after the onset of exercise) and 5 features (corresponding to CGM reading, insulin, carbohydrate, heart rate and step count in order). We will denote the first feature (CGM reading) as $y$ and the rest 4 features as $x$. We will use subscript to indicate discrete time steps and superscript to indicate feature indices. For example, $x_{1}^{2}$ is the 2nd feature of $x$ (carbohydrate) at discrete time step $t=1$. To better indicate past and future, we set the 7th last entry to be $t=0$ and define 1 step of time to be 5 minutes. Our goal is to predict the CGM trace during the first 30 minutes of exercise corresponding to the expected output $y_{1:6}\in\mathbb{R}^{6}$ and therefore we set the number of prediction steps $q$ to be 6 for all models. We further split the given time series into historical context $p=(y,x)_{-47:-1}\in\mathbb{R}^{47\times 5}$, starting glucose $y_{0}\in\mathbb{R}$, inputs during exercise $x_{0:5}\in\mathbb{R}^{6\times 4}$ (six inputs that are recorded 1 time step ahead of the expected outputs). We use $\hat{y}\in\mathbb{R}^{6}$ to indicate the CGM trace predicted by models, $s$ for modeled states and $z$ for latent states. $h,c$ for the final hidden state and cell state of the LSTM initial condition learner. For ease of computation and without loss of generality, we set the $\Delta t$ term in forward-Euler style discretization to be $1$ for all relevant models, and thus we omit it in the equations.
Learning Rate, Training Epochs and Optimizer
Also for all models, we set the default learning rate to be $2\cdot 10^{-3}$ and use the Adam optimizer (Kingma and Ba, 2014). Unless otherwise specified in the model description, we train for 100 epochs and pick the best epoch with a validation set. For each model we tested this learning rate and training epoch on an initial try-it-out run and found them to be robust for all the models (in most cases training error converges within first 50 epochs). While we understand that the best practice is to always tune learning rates, here we made the decision not to tune it in an attempt to reduce the dimension of the hyperparameter search space, which is already very large for some models. Since Adam already adjusts learning rate for each parameter in the model, we did not see significant variation in model performance with other learning rates.
@@
One may notice that for the hybrid models, we place 0 dropout rate on the MLPs. This is because we are already doing regularization via epoch selection based on validation set, and our trial runs indicate that there is no need for additional regularization (adding any non-zero dropout for hybrid models worsens performance). This is not true for black-box models, for which we still tune the dropout rate.
@@
We use grid search to tune hyperparameters. When choosing the grid, we restrict the search space to areas where the models have less than 25000 parameters and we also try to limit the number of grid points to around 10. This is to make sure the computational cost of the experiments is capped at a reasonable level for small data sets and individual users.
@@ E.1 UVA
Inside the algorithm, the LSTM network has 2 layers and 21 hidden dimensions (same as the UVA/Padova model), and we set the first state of the estimated initial condition (represented by hidden state $h$) to be the true initial value of CGM $y_{0}$ to keep consistency with the assumption that $s^{1}$ represents glucose. The UVA/Padova Simulator has 52 trainable parameters $\beta\in\mathbb{R}^{52}$. We do not tune hyperparameters for the UVA model as there are not meaningful hyperparameters to tune.
@@ E.2 Reduced UVA Latent Parameter Learning
Note in the above implementation, the latent dynamics of $z$ only depends on itself and inputs that are not used by the UVA S2013 model (the 3rd and 4th feature, which correspond to heart rate and step count). We made this choice to keep consistency with the implementation in Miller et al. (2020). The LSTM has 2 layers and $d$ hidden dimension (since this time the cell states are used to initialize latent state $z$, whose dimension needs to be tuned) and MLP2 serves to map $d$ dimensional $h_{0}$ to a 8-dimensional initial condition vector. All MLPs have $n$ hidden layers and $m$ hidden units with dropout 00 and activation ReLu, we tune these hyperpameters with grid search on the following grid:
@@ E.3 Reduced UVA Latent Parameter and State Closure Learning
  Input: number of prediction steps $q=6$, historical context $p$, starting glucose $y_{0}$, exogenous inputs $x_{0:5}$, the reduced UVA mechanistic ODEs $m_{\text{RUVA}}$ the adjacency matrices (as defined in Section 3) of reduced UVA: $A_{s},A_{x}$, closure switch constant $c$, $\Delta t=1$
@@
     $s_{i+1}=s_{i}+\Delta t\cdot m_{\text{RUVA}}(s_{i},x_{i};\beta_{i}=\text{MLP2}(z_{i}))+c\cdot\text{MLPs}(s_{i},x_{i};A_{s},A_{x})$
@@
The training routine of models adopting state closure is more complicated. We first set $c=0$ and just train the reduced UVA latent parameter model and the initial condition learner LSTM for 100 epochs (the closure part is masked out by $c$). Then we freeze the weights of the reduced UVA latent parameter model and the LSTM, set $c=1$ and train the closure neural network on the residuals for 50 more epochs. This trick makes sure the closure network is truly learning the residuals as intended and not subsuming the mechanistic model.
@@ E.5 BNODE
For BNODE and the subsequent black-box models, we do not provide design rationales behind them as it is better to refer to their original author for such details. Instead, we simply describe our how we implemented them.
@@ E.6 TCN
The TCN model is taken directly from the code repository posted on https://github.com/locuslab/TCN/blob/master/TCN/tcn.py, with input size set to 5, output size set to 6, number of channels set to a list of $n$ copies of $m$, kernel size set to $l$ and dropout set to $a$. We tune these hyperpameters with grid search on the following grid:
@@ E.7 LSTM
  $seq_{out}h_{q},c_{q}=\text{Decoder LSTM}(x_{0:5})$
@@ E.8 Transformer
  Input: number of prediction steps $q=6$, historical context $p$, starting glucose $y_{0}$, exogenous inputs $x_{0:5}$, true output $y_{1:5}$ (needed during training)
@@
We use the transformer model provided by the pytorch nn class, and its hyperparameters are set as follows: d_model set to $n$, number of encoder layers and number of decoder layers are both set to 3, the dim_feedforward is set to $m$ and dropout is set to $a$. We tune the hyperparameters with the following grid:
$n=\{8,16\}\times m=\{32,64\}\times a=\{0,0.1,0.2\}$
@@ E.9 S4D
  $\hat{y}=\text{Linear}(seq_{out})_{-6:}$
@@
We take the S4D model directly from the following github repository https://github.com/thjashin/multires-conv/blob/main/layers/s4d.py, and its hyperparameters are set as: d_model set to $53$, d_state set to $m$, dropout set to $a$.We tune the hyperparameters with the following grid:
$m=\{32,64,128\}\times a=\{0,0.1,0.2\}$

### v2

@@ D.1 T1DEXI Exercise Instances
    CGM readings. CGM readings are taken every 5 minutes for all patients in T1DEXI.
@@
    Insulin injection. If the patient injected insulin at a given time in the T1DEXI dataset, we add that amount of insulin to the corresponding 5 minute interval to create a bolus insulin injection time series. We add basal insulin to the bolus insulin time series to produce the (total) insulin time series used in our experiments.
@@
    Carbohydrate intake. Suppose the patient consumed carbohydrates at a given time in the T1DEXI dataset. We assume a constant meal consumption rate of 45 grams per minute, and then compute the average consumption rate over each 5 minute interval to create a carbohydrate consumption time series.
@@
    Heart rate. We average the heart rate over each 5 minute interval to obtain a heart rate time series.
  5. 5.
    Step count. Similarly, we average step counts over each 5 minute interval to obtain a step count time series.
We only consider exercise instances for which:
  1. 1.
    all 54 CGM readings were available;
  2. 2.
    the patient logged carbohydrates during the instance;
  3. 3.
    the exercise duration was at least 30 minutes;
  4. 4.
    it was either the first or second exercise instance for that patient; and
  5. 5.
    HR was consistently recorded every 10 seconds without any missingness throughout the 54 5-min intervals.
After this process we end up with 143 exercise instances (13 patients with 1 exercise instance, and 65 patients with 2 exercise instances).
Figure 12 depicts the raw and processed data for a single exercise time-window for a single patient. The first row shows a complete timeseries of CGM, while the second row shows the recorded changes in basal insulin rates. While insulin pump data only includes the change points (red dots in the second plot), the devices function by executing the new constant rate (blue line) until a new basal rate is set (a subsequent red dot). The third plot shows bolus insulin recordings (red dots), which are sometimes thought of as being delivered instantaneously. However, insulin pumps in fact administer these doses by delivering insulin at a constant rate over a suitable short time window (typically $<5$ min), such that the sum total of disbursed bolus (integral of the blue curve in the third plot) corresponds to the recorded dose (red dot). In the case of insulin pumps, both the basal and bolus insulins are fundamentally the same drug and are only distinguishable by their administration pattern; for this reason, we convert all insulin administration to $U/min$ and sum the basal and bolus, producing the black curve in the fourth figure.
The fifth plot shows the amount of carbohydrates ($g$) that were recorded by the participant (either as a meal or as part of the clinical study as “rescue” carbohydrates explicitly intended to mitigate hypoglycemia). We assume that meals are consumed at a constant $45g/min$ and plot the corresponding carbohydrate consumption rate in the sixth plot (blue); we then align this impulse function with the 5 minute grid defined by CGM data (black). Finally, the last two figures show heart rate and step counts, respectively. In both cases, we average the raw data (red) over each 5 minute window, and perform piecewise-constant interpolation of these values (blue). Observe that step count and heart rate begin to rise at $t=-60$; the exercise lasts for approximately 1 hour, then the step counts and heart rates drop for $t>0$, once exercise is complete.
Figure 12: We show an example of the raw and processed data from a single patient window. Here, time 00 represents the end of exercise, and we plot all features for the 4 hours prior to exercise termination, as well as the subsequent 30 minutes. In all plots, red dots indicate raw measurements, blue lines represent our clinically-informed representations and interpolations of the raw data, and the black curves (when relevant) represent subsequent mapping of the interpolated data to 5 minute intervals.
D.2 Intervention Sets
For each exercise time series, we first create an intervention set containing three copies of the original time series. And then we uniformly randomly (simulated by the numpy default random number generator with seed 2024) apply one of the four following intervention sets to it:
  1. 1.
    Add 0/50/100 grams of carbohydrate to the three copies at the end of exercise (time step $t=0$);
  2. 2.
    Add 0/2.5/5 units of insulin to the three copies from the end of exercise onwards (time steps $t=0,1,\ldots,6$);
  3. 3.
    Add nothing/50 grams of carbohydrate/10 units of insulin to the three copies at the end of exercise (time step $t=0$);
  4. 4.
    Change the recorded heart rate in the last 7 time steps $t=0,1,\ldots,6$ to typical medium intensity aerobic training (80,90,100,110,120,130,120)/ typical interval training (80,170,80,170,80,170,80)/typical high intensity resistance training (160,170,180,170,160,180,160).
For each exercise instance this leads to four time series (each consisting of 5 features and 54 time steps): the original observed time series, and three copies corresponding to the intervention set.
In addition, we compute the class label for each intervention set as the index of the intervention that leads to the highest mean glucose: 100 grams of carbohydrates for category 1; zero insulin for category 2; 50 grams of carbohydrate for category 3; and the highest heart rate level for category 4.
@@ Appendix E Model Implementations and Hyperparameters
For each model mentioned in the experiment section, here we offer a detailed description of the corresponding computational method and the hyperparameters used. Throughout this section we are given an exercise time series of 54 time steps (corresponding to 54 5 minute intervals that made up the time window starting from 4 hours prior to exercise termination, and ending at 30 minutes after exercise termination) and 5 features (corresponding to CGM reading, insulin, carbohydrate, heart rate, and step count, in that order). We denote the first feature (CGM reading) as $y$, and the other 4 features as $x$. We use a subscript to indicate discrete time steps and superscript to indicate feature indices. For example, $x_{1}^{2}$ is the 2nd feature of $x$ (carbohydrate) at discrete time step $t=1$.
Our goal is to predict the CGM trace during the first 30 minutes following exercise completion corresponding to the output $y_{1:6}\in\mathbb{R}^{6}$ and therefore we set the number of prediction steps $q$ to be 6 for all models. We further split the given time series into historical context $p=(y,x)_{-47:-1}\in\mathbb{R}^{47\times 5}$, starting glucose $y_{0}\in\mathbb{R}$, inputs during exercise $x_{0:5}\in\mathbb{R}^{6\times 4}$ (six inputs that are recorded 1 time step ahead of the expected outputs). We use $\hat{y}\in\mathbb{R}^{6}$ to indicate the CGM trace predicted by models, $s$ for modeled states and $z$ for latent states. $h,c$ for the final hidden state and cell state of the LSTM initial condition learner. For ease of computation and without loss of generality, we set the $\Delta t$ term in forward-Euler style discretization to be $1$ for all relevant models, and thus we omit it in the equations.
Learning Rate, Initialization and Optimizer
For all experiments, we use the Adam optimizer (Kingma & Ba, 2015) to perform gradient descent.
Learning rates and weight initialization are set with model stability as a consideration. In particular, in all experiments using T1DEXI data, we set the default learning rate to be $2\times 10^{-3}$ for all models except UVA, for which we use a larger learning rate of $10^{-1}$. For all models except UVA, we initialize model weights with the PyTorch default setting if they are part of a pre-defined PyTorch model class or standard normal if they are custom weights. For UVA we initialize model weights from a normal distribution with mean zero and variance $1/400$. UVA is treated differently because ODE-based mechanistic models are notoriously highly sensitive to parameter initialization, and it is both important to initialize the parameters to be small in magnitude, and to make sure they do not get too close to zero during training.
For synthetic experiments, we set the learning rate of the purely mechanistic model to $5\times 10^{-1}$; to $1\times 10^{-2}$ for LP and LPSC; and to $2\times 10^{-3}$ for the remainder of the models.
Training Epochs
Unless otherwise specified in the model description (LPSC), in T1DEXI experiments, we train for 100 epochs and pick the epoch with best validation loss. In synthetic experiments, we train for 50 epochs and pick the epoch with best validation loss.
@@
For the hybrid models, we use 0 dropout rate. This is because we are already doing regularization via early stopping, and our trial runs indicate that there is no need for additional dropout. This is not true for black-box models, for which we still tune the dropout rate.
@@
We use grid search to tune hyperparameters. When choosing the grid, we restrict the search space to areas where the models have less than 20000 parameters and we also try to limit the number of grid points to around 10. This is to make sure the computational cost of the experiments is capped at a reasonable level for small data sets and individual users. The grid used for each model in each experiment will be provided below together with model descriptions.
@@ E.1 UVA
T1DEXI Experiments
The LSTM network has 2 layers and 21 hidden dimensions (same as the UVA/Padova model), and we set the first state of the estimated initial condition (represented by hidden state $h$) to be the true initial value of CGM $y_{0}$ to keep consistency with the assumption that $s^{1}$ represents glucose. The UVA/Padova Simulator has 53 trainable parameters $\beta\in\mathbb{R}^{53}$. We do not tune hyperparameters for the UVA model.
Synthetic Experiments
In the synthetic setting, we assume to know the ground truth model. And therefore we remove the LSTM initial condition encoder and instead set $s_{0}=y_{0}$ since in the synthetic setting, there is only one state. Same as T1DEXI, we do not tune hyperparameters.
@@ E.2 Reduced UVA Latent Parameter Learning
T1DEXI Experiments
In our implementation, the latent dynamics of $z$ only depends on itself and inputs that are not used by the UVA S2013 model (the 3rd and 4th feature, which correspond to heart rate and step count). We made this choice to keep consistency with the implementation in Miller et al. (2020). The LSTM has 2 layers and $d$ hidden dimension (since this time the cell states are used to initialize latent state $z$, whose dimension needs to be tuned) and MLP2 serves to map $d$ dimensional $h_{0}$ to a 8-dimensional initial condition vector. All MLPs have $n$ hidden layers and $m$ hidden units with dropout 00 and activation ReLu, we tune these hyperpameters with grid search on the following grid:
@@
Synthetic Experiments
Similar to the purely mechanistic model, in the synthetic setting we directly set $s_{0}=y_{0}$ and therefore do not use $h$ or MLP1. We still need the 2-layer LSTM to generate $z_{0}$. The search grid for synthetic experiments is:
$n=\{2,3\}\times m=\{16,32\}\times d=\{2,4\}$
@@ E.3 Reduced UVA Latent Parameter and State Closure Learning
  Input: number of prediction steps $q=6$, historical context $p$, starting glucose $y_{0}$, exogenous inputs $x_{0:5}$, the reduced UVA mechanistic ODEs $m_{\text{RUVA}}$ the adjacency matrices (as defined in Section 3) of reduced UVA: $A_{s},A_{x}$, closure switch constant $w$, $\Delta t=1$
@@
     $s_{i+1}=s_{i}+\Delta t\cdot m_{\text{RUVA}}(s_{i},x_{i};\beta_{i}=\text{MLP2}(z_{i}))+w\cdot\text{MLPs}(s_{i},x_{i};A_{s},A_{x})$
@@
Synthetic Experiments
We again set $s_{0}=y_{0}$ and do not use $h$ or MLP1. The hyperparameter grid is:
$n=\{2,3\}\times m=\{16,32\}\times d=\{8,16\}$
Special Training Routine
The training routine of models adopting state closure is more complicated. We first set $w=0$ and just train the reduced UVA latent parameter model and the initial condition learner LSTM for 100(T1DEXI)/50(Synthetic) epochs (the closure part is masked out by $w$). Then we freeze the weights of the reduced UVA latent parameter model and the LSTM, set $w=1$ and train the closure neural network on the residuals for 50 more epochs. This trick makes sure the closure network is truly learning the residuals as intended and not subsuming the mechanistic model.
@@ E.4 Mechanistic Neural ODE
Synthetic Experiments
We set $s_{0}=y_{0}$ and do not use $h$ or the LSTM (note that our code implementation still has the LSTM component for the sake of consistency but its output is actually never used by MNODE.) The search grid is:
$n=\{2,3\}\times m=\{16,32\}.$
@@ E.5 BNODE
For BNODE and the subsequent black-box models, we point the reader to implementations referenced in the associated citations in the main paper. Here we describe our implementation.
@@
Synthetic Experiments
We keep all the settings the same (note we can no longer assume there is only 1 state as in the black-box regime we have zero domain knowledge) as the T1DEXI experiments except that the search grid for hyperparameters is now:
$d=\{2,3\}\times n=\{2,3\}\times m=\{64\}\times a=\{0,0.2\}.$
@@ E.6 TCN
T1DEXI Experiments
The TCN model is taken directly from the code repository posted on https://github.com/locuslab/TCN/blob/master/TCN/tcn.py, with input size set to 5, number of channels set to a list of $n$ copies of $m$, kernel size set to $l$ and dropout set to $a$. We tune these hyperpameters with grid search on the following grid:
@@
Synthetic Experiments
We keep all the settings the same as the T1DEXI experiments except that the input size is set to 3 and the search grid for hyperparameters is now:
$n=\{2,3\}\times m=\{16,32\}\times l=\{2,3,4\}\times a=\{0,0.2\}.$
@@ E.7 LSTM
  $seq_{out},h_{q},c_{q}=\text{Decoder LSTM}(x_{0:5})$
@@ E.8 Transformer
  Input: number of prediction steps $q=6$, historical context $p$, starting glucose $y_{0}$, exogenous inputs $x_{0:q-1}$, true output $y_{1:q-1}$ (needed during training)
@@
T1DEXI Experiments
We use the transformer model provided by the pytorch nn class, and its hyperparameters are set as follows: d_model set to $d$, number of encoder layers set to $l_{1}$, number of decoder layers set to $l_{2}$, the dim_feedforward is set to $m$ and dropout is set to $a$. We tune the hyperparameters with the following grid:
$d=\{4,8\}\times l_{1}=\{2,3\}\times l_{2}=\{2,3\}\times m=\{32,64\}\times a=\{0,0.1\}$
Synthetic Experiments
We keep all experiment settings the same except setting $q=10$ and the hyperparameter search grid to be:
$d=\{4,8\}\times l_{1}=\{2\}\times l_{2}=\{2\}\times m=\{32,64\}\times a=\{0,0.2\}$
@@ E.9 S4D
  Input: number of prediction steps $q=6$, historical context $p$, starting glucose $y_{0}$, exogenous inputs $x_{0:q-1}$
@@
  $seq_{out}=\text{Transpose}(Seq_{out},1,2)$
  $\hat{y}=\text{Linear}(seq_{out})_{-q:}$
@@
T1DEXI Experiments
We take the S4D model directly from the following github repository https://github.com/thjashin/multires-conv/blob/main/layers/s4d.py, and its hyperparameters are set as: d_model set to $d$, d_state set to $m$, dropout set to $a$.We tune the hyperparameters with the following grid:
$d=\{4,6,8\}\times\{m=\{32,64\}\times a=\{0,0.1,0.2\}$
Synthetic Experiments
We set $q=10$ and the hyperparameter search grid to be:
$d=\{3,4,5\}\times\{m=\{32,64\}\times a=\{0,0.2\}$
@@ Appendix F Repeated Nested Cross Validation
Appendix G Additional Experiments
In this section we provide details of additional experiments that may be of interest to readers. We carried out three sets of additional experiments: (1) we apply novel interventions to the test set that are unseen in the training set; (2) we corrupt a small portion of training ranking information; and (3) we test hybrid models that are based on full (un-reduced) mechanistic models.
G.1 Different Intervention Sets in Training and Test Sets
In this experiment, the training intervention sets consist of only category (1) and (2) described in Section 5.2 (Intervention Sets). Crucially, note that our training intervention sets do not include interventions where both insulin and carbohydrates can vary, as in category (3) in Section 5.2; this ensures that our training is only teaching the models about dependence on insulin alone and carbs alone. On the other hand, the test intervention set contains only the following category: add 45 grams of carbohydrate at the end of exercise, and at the same time add 2.25/3.00/4.50 units of insulin. These interventions correspond to a insulin-carb ratio of 1:20/1:15/1:10 respectively. We keep other experimental settings the same as the main T1DEXI experiments.
Results
The results of the experiments are summarized in Figure 13. The results illustrate that for moderate values of $\alpha$ that balance the prediction loss with the causal loss, the causal error rate of most models improves even in this setting where the test intervention set contains interventions not seen in the training data. Further, we see that in general, the hybrid models offer more rapid improvement in causal loss as $\alpha$ increases, with virtually no increase in predictive loss. BNODE appears to offer comparable performance to MNODE, but with generally higher variance in its causal loss.
Figure 13: As in Fig. 4, but for T1DEXI data and using a full UVA/Padova mechanistic model baseline and hybridizations of the reduced UVA/Padova simulator for LP, LPSC, and MNODE. In these experiments, the training data includes only interventions on either insulin or carb intake, but not both together; however, the test data includes interventions on the insulin-carb ratio.
G.2 Corrupted Ranking Knowledge in Training Set
In this experiment, the ground truth ranking of each training intervention set in $\mathcal{I}^{*}$ has a probability of $\rho\in\{0.05,0.10,0.20\}$ of being corrupted by circularly shifting the ranking to the right by 1 position. We carry out these simulations only for MNODE as a representative example. We keep the testing ground truth intact and other experimental settings the same as the main T1DEXI experiments.
Results
The results of the experiments are summarized in Figure 14. We see that at moderate values of $\alpha$, the performance on causal tasks degrades with increasing levels of corruption. Notably, we also see that when the corruption rate is high, high $\alpha$ (high weight on the causal loss) can actually deteriorate causal performance, suggesting that the model is erroneously overemphasizing rankings from the corrupted data (introducing bias). Regardless, even when the corruption rate is high, choosing any $\alpha>0$ (i.e., using a hybrid loss) lowers the causal error rate relative to a pure predictive loss ($\alpha=0$).
Figure 14: As in Fig. 4, but for T1DEXI data with MNODE, using a hybridization of the reduced UVA/Padova simulator. In these experiments, a fraction $\rho$ of the training data is corrupted.
G.3 Hybrid Models Based on Full UVA/Padova
In the main paper, we work with a reduced UVA/Padova model. For completeness, we also implemented the latent parameter (LP) model and mechanistic neural ODE (MNODE) model using the full UVA/Padova model described in Appendix A. Their implementations are exactly the same as described in Appendix E except that they use model equations/adjacency matrices of the full UVA/Padova model, and their hyperparameters are tuned on slightly different grids to count for the drastic increase in the number of mechanistic states and parameters. The search grid for hyperparameters of each full hybrid model is given below:
  1. 1.
    LP:
    $n=\{2,3\}\times m=\{16,24,32\}\times d=\{20,24,28\}$
  2. 2.
    MNODE:
    $n=\{2,3\}\times m=\{16,32\}.$
Note that we did not include latent parameter learning with state closure (LPSC) in this set of simulations, as its full version requires an unreasonably large number of both states and model parameters when applied to the full UVA/Padova model.
Results
The results of the experiments are summarized in Figure 15. We see that in general, the hybrid models based on the reduced UVA/Padova simulator offer generalization performance on par with those of hybrid models based on the full simulator. One of the primary benefits of building off of a reduced mechanistic model is that the resulting hybrid model has significantly reduced training time. These results suggest model reduction is a valuable (but not necessary) implementation step in applying H2NCM in practice.
Figure 15: As in Fig. 4, but for T1DEXI data and comparing LP and MNODE hybrid models based on both the full and reduced UVA/Padova simulator as the mechanistic model.